If FTL is 56 N, FW is 100 N, and the box is not moving, what is FTR?
Answer #1

44N

Beacuse Upward Forces should balance downward Forces.

FTL + FTR =FW

56+FTR =100

FTR = 100-56

= 44
Answer #2

FTL + FTR = FW (since box is stationary)

FTR = FW - FTL = 100 - 56 = 44N
Answer #3

BOX is not moving so forces acting on boxes are in equlibrium or in balance so just balance all the forces


Fw = Ftl+Ftr

this in balance

so

100 N = Ftr +56N

Ftr = 44N